Explain why solving a system of equation graphically can be difficult. Use y=2x-5 and y=-x+3.
Answer by ewatrrr(24785)

 
Hi
y=2x-5 ||(x,y) ordered pairs (0,-5) and (5/2,0) on this Line. Plot points & connect with Line
y=-x+3. || (x,y) ordered pairs (0,3) and (3,0) on this Line. Plot points & connect with Line
Explain why solving a system of equation graphically can be difficult:
fractional values may be involved in the (x,y) ordered pair of the Intersection
2x-5 = -x+3 , x = 8/3 and y = 1/3 , Pt(8/3,1/3) point of Intersection